### v2

# Ordering in statistical systems on the way to the thermodynamic limit

Abstract
It is well known that the mathematically accurate description of ordering and related symmetry breaking in statistical systems requires to consider the thermodynamic limit. But the order does not appear from nowhere, and yet before the thermodynamic limit is reached, there should exist some kind of preordering that appears and grows in the process of increasing the system size. The quantitative description of growing order, under the growing system size, is developed by introducing the notion of order indices. The rigorous proof of the phase transition existence is a separate difficult problem that is not the topic of the present paper. We illustrate the approach resorting to several models in the mean-field approximation, which makes it possible to demonstrate the notion of order indices for finite systems in a clear way. We show how the order grows on the way to the thermodynamic limit for Bose-Einstein condensation, arising superconductivity, magnetization, and crystallization phenomena.
Keywords: Statistical systems, Thermodynamic limit, Order indices, Reduced density operators, Correlation operators
1 Introduction
Different thermodynamic phases are usually characterized by order parameters related to the specific symmetries of the phases. The order parameters, as is well known, become meaningful only in the thermodynamic limit, when the number of particles in the system and its volume tend to infinity. Only in this limit, it is possible to give a mathematically correct description of spontaneous symmetry breaking and the related phase transitions. For example, it is straightforward to show that, if a spin Hamiltonian possesses rotational or reversal symmetry, the average spin, hence magnetization, is identically zero for a finite system. Similarly, if a Hamiltonian enjoys global gauge symmetry, Bose-Einstein condensation cannot appear in a finite system. In the same way, when a Hamiltonian is translationally invariant, equilibrium crystalline state cannot arise in a finite system. Detailed discussion and mathematically accurate description of these facts can be found in the books [1, 2, 3, 4, 5, 6].
In order to describe phase transitions and spontaneous symmetry breaking in mathematically accurate way, Bogolubov [1, 2, 4] developed the method of quasi-averages, where the Hamiltonian symmetry is broken by an additional term that is removed after the thermodynamic limit. It is important to stress that the thermodynamic limit goes first, and the symmetry-breaking term is removed only after that. In principle, it is possible to define the symmetry-breaking term so that it would go to zero almost simultaneously with, but anyway a bit slower than the thermodynamic limit [7]. In any case, the thermodynamic limit is compulsory for the mathematically rigorous definition of order parameters.
A principal question that has remained unclear is wether the ordering could happen before the thermodynamic limit, when the system is already sufficiently large, although not yet infinite, and how the order could grow under the increasing system size until the thermodynamic limit. It is the aim of the present paper to suggest a mathematical description for the process of order growth under the increasing system volume up to the thermodynamic limit.
In order to avoid misunderstanding, let us specify what is the main picture we keep in mind and what are the secondary details that can be neglected and that we will not touch. We consider a system of NN particles in volume $V$, hence density $\rho=N/V$, under fixed thermodynamic and external parameters. According to the ideology of the thermodynamic limiting procedure, we then consider the system of a larger number of the same kind of particles in a larger volume, such that the density remains constant and all thermodynamic parameters are the same, being fixed. In that way, we consider not a single system, with increasing volume, but a sequence of systems with increasing sizes keeping the same fixed density and thermodynamic parameters. This is exactly the procedure accepted in the thermodynamic limit.
We consider statistical systems with large numbers of particles $N\gg 1$. Because of this inequality, although the number NN can be finite, but, due to its large value, boundary effects can be neglected, especially when $N\rightarrow\infty$. Therefore we do not need to resort to the thermodynamics of small systems of the size of single molecules [8]. Since, we increase the system size explicitly, we do not need to involve finite-size scaling [9] that is used in numerical calculations, when one tries to extrapolate to an infinite system size the measurements accomplished for small systems.
Of course, sometimes, when considering finite systems of intermediate size, one needs to take account of surface effects. For instance, the properties of finite clusters of magnetic materials can depend on surface properties, inducing, e.g., magnetic anisotropy [10].
In certain situations, finite-size effects can play a role. For example, even versus odd number of fermions can play a role for atomic nuclei and metallic clusters. Shape and orientation transitions can occur for small systems in external fields or under rotation [11]. Anisotropy in geometry is important for studying one- and two-dimensional limiting cases [12]. However, we keep in mind the standard understanding of large systems as those for which boundary effects, although could yield some small corrections, but are not important for the considered phenomena, so that to a good approximation can be neglected when $N\gg 1$ [13]. We assume that there are no external fields imposing anisotropy throughout the whole system.
We do not aim of studying the behavior of thermodynamic characteristics in the close vicinity of critical points. Therefore we do not consider the semi-phenomenological droplet approximation [14, 15, 16] for describing the critical region. In droplet models one considers large systems as composed of many small droplets, with the energy and entropy of each of them described by the so-called leptodermous expansions in powers of the droplet volume, with phenomenological coefficients. In that way, a droplet approximation treats a large single system as being composed of numerous small droplets. This picture does not have much relation to our aim of considering a sequence of systems whose volume increases tending to the thermodynamic limit.
Each physical system, under the fixed density and other thermodynamic parameters, is assumed to be equilibrium. In an infinite system, this can be an absolute equilibrium while in finite large systems it is rather a metastable equilibrium. A mathematically rigorous characterization of metastable states can be found in the review by Sewell [17]. Being interested in equilibrium states, we do not touch nonequilibrium phenomena related to the nucleation and temporal formation of a new thermodynamic phase or a new structure, starting from a state of metastability [18, 19, 20, 21].
Thus, we consider a sequence of equilibrium statistical systems differing from each other only by the size, that is by the number of particles and system volume, with fixed density and all thermodynamic and external parameters. Each system consists of a large number of particles, such that surface effects can be neglected. Our aim is to develop a quantitative characterization of the evolving order under the increasing system size on the way to the thermodynamic limit.
The guiding idea for developing a mathematical approach that could describe the growing order under the increasing system size can be qualitatively characterized as follows. Order, emerging in a system, depends on the relation between an effective correlation length and the system size. The correlation length is defined by the corresponding correlation function. For example, a particular important case of correlation functions is represented by reduced density matrices [22]. The latter can be treated as matrix elements of the reduced density operators $\hat{\rho}_{n}$, where $n=1,2,\ldots$. The norm of a density operator, $\hat{\rho}_{n}||$, is connected with its largest eigenvalue, which, in turn, is related to the effective correlation length. The trace of the density operator is a function of the number of particles NN. For instance, $\rm{Tr}\rho_{1}=N$. Then the exponent $\omega$, characterizing the relation between the norm and trace of a density operator, describes the dependence of the density operator spectrum on the number of particles, $\hat{\rho}_{1}||=N^{\omega}$. The parameter $\omega$ is an order index showing how quickly the order grows with the increase of the system size. When the order index is small, $\omega\ll 1$, this implies that the correlation length is much shorter that the system effective length, hence there is no long-range order. But when the order index is close to one, then the correlation length becomes comparable to the system effective size, which means that long-range order is getting established in the system.
It is the aim of the present paper to formulate these ideas, employing reduced density matrices for finite systems, in a rigorous quantitative way, to extend the consideration to general correlation operators, and to illustrate the approach by several concrete cases of arising long-range order in the process of growing system size tending from finite NN to the thermodynamic limit. We consider the cases of Bose-Einstein condensation, superconducting transition, magnetic transition, and crystal-liquid transition, which, to make the ideas transparent, are treated in mean-field approximation.
Throughout the paper, we use the system of units, where the Planck constant $\hbar$ and the Boltzmann constant $k_{B}$ are set to one.
2 Operator order indices
In the thermodynamic limit, the level of order of macroscopic systems can be characterized by order indices of reduced density matrices [23, 24, 25]. Here we extend the notion of order indices of reduced density matrices to finite statistical systems and generalize this notion to arbitrary correlation operators, showing that this notion provides a direct tool for measuring the level of order increasing together with the system size on the way to the thermodynamic limit.
Order indices can be introduced for any trace-class operator, for which
[…]
In what follows, we shall deal with semi-positive operators. Let a trace-class semi-positive operator $\hat{A}$ be defined on a Hilbert space $\mathcal{H}$. The operator order index is
[…]
Here the trace is taken over a Hilbert space $\mathcal{H}$ and the logarithm can be taken with respect to any base. Below, we prefer to use the natural logarithm. This index shows how the operator norm depends on the operator trace,
$\hat{A}\;||\;=\;|\;{\rm Tr}\;\hat{A}\;|^{\;\omega(\hat{A})}\;$.
For semi-positive operators,
$\hat{A}\;||\;\leq\;{\rm Tr}\;\hat{A}\qquad(\hat{A}\geq 0)\;,$ (2)
then the order index is limited from above by one,
$\omega(\hat{A})\;\leq\;1\qquad(\hat{A}\geq 0)\;.$ (3)
From below, the index is not necessarily limited. There is order in the operator, when $\omega(\hat{A})>0$, while there is no order, if $\omega(\hat{A})\leq 0$. Increasing order index implies the growth of order.
In general, the norm can be introduced in different ways. The operator norm
[…]
where
$\hat{A}\varphi\;||\;=\;\sqrt{\langle\;\varphi\;|\;\hat{A}^{+}\hat{A}\;|\;\varphi\;\rangle}\;,$
seems to be the most appropriate for the following applications. For Hermitian operators, the operator norm reduces to the Hermitian norm
[…]
This norm will be used throughout the paper, since we shall deal with Hermitian operators.
Note that semi-positive trace-class operators are bounded, since
[…]
Hence the order indices are well defined for this class of operators.
3 Reduced density operators
The notion of order indices, introduced above for arbitrary operators, can describe the ordering processes in physical systems when this notion is applied to the operators characterizing mutual correlations in these systems. The detailed information on the properties of statistical systems is contained in the reduced density matrices [22]. These matrices can be treated as matrix elements of reduced density operators.
Thus, the first-order reduced density matrix, defined by the statistical average
$\rho(x,x^{\prime})\;=\;{\rm Tr}\;\psi(x)\;\hat{\rho}\;\psi^{\dagger}(x^{\prime})\;=\;\langle\;\psi^{\dagger}(x^{\prime})\;\psi(x)\;\rangle\;,$ (7)
where $x$ is a set of physical variables, like spatial coordinates, spin etc, can be treated as a matrix element of the first-order reduced density operator
$\hat{\rho}_{1}\;=\;[\;\rho(x,x^{\prime})\;]\;.$ (8)
This operator is defined on the Hilbert space that is the closed linear envelope
${\cal H}_{1}\;=\;\overline{{\cal L}}\{\;|\;k\;\rangle\;\}$ (9)
over the basis formed by natural orbitals
$|\;k\;\rangle\;=\;|\;\varphi_{k}\;\rangle\;=\;[\;\varphi_{k}(x)\;]$ (10)
reflecting the system properties [22], with $k$ being a set of quantum numbers labeling the orbitals.
The norm of the first-order density operator is
$\hat{\rho}_{1}\;||\;=\;\sup_{k}\;N_{k}\;,$ (11)
where
$N_{k}\;\equiv\;\langle\;k\;|\;\hat{\rho}_{1}\;|\;k\;\rangle\;=\;\int\varphi_{k}^{*}(x)\;\rho(x,x^{\prime})\;\varphi_{k}(x^{\prime})\;dxdx^{\prime}$ (12)
plays the role of the occupation number associated with the state labeled by $k$. The trace of operator (8) over the space (9) is
${\rm Tr}\;\hat{\rho}_{1}\;=\;\sum_{k}\langle\;k\;|\;\hat{\rho}_{1}\;|\;k\;\rangle\;=\;\int\rho(x,x)\;dx\;=\;N\;.$ (13)
Then the order index for this operator, as defined in Ref. (1), is
$\omega(\hat{\rho}_{1})\;=\;\frac{\log\;\sup_{k}N_{k}}{\log N}\;.$ (14)
As is seen, the order index shows how quickly the largest occupation number increases under the growing system size comprising NN particles,
$\sup_{k}N_{k}\;=\;N^{\omega(\hat{\rho}_{1})}\;.$
Similarly, the second-order reduced density operator
$\hat{\rho}_{2}\;=\;[\;\rho_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;]$ (15)
is a matrix of the elements
$\rho_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;=\;\langle\;\psi^{\dagger}(x_{2}^{\prime})\;\psi^{\dagger}(x_{1}^{\prime})\;\psi(x_{1})\psi(x_{2})\;\rangle$ (16)
enjoying the property
$\rho_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;=\;\rho_{2}(x_{2},x_{1},x_{2}^{\prime},x_{1}^{\prime})\;.$ (17)
This operator acts on the Hilbert space
${\cal H}^{2}\;=\;{\cal H}_{1}\bigotimes{\cal H}_{2}\;=\;\overline{{\cal L}}\{\;|\;kp\;\rangle\;\}$ (18)
that is the closed linear envelope over the basis formed by the functions
$|\;kp\;\rangle\;=\;|\;\varphi_{k}\;\rangle\;\bigotimes\;|\;\varphi_{p}\;\rangle\;=\;[\;\varphi_{k}(x)\;]\;\bigotimes\;[\;\varphi_{p}(x)\;]\;.$ (19)
The norm of the second-order operator (15) reads as
$\hat{\rho}_{2}\;||\;=\;\sup_{kp}\;N_{kp}\;,$ (20)
where
$N_{kp}\;=\;\langle\;kp\;|\;\hat{\rho}_{2}\;|\;kp\;\rangle\;=$
$\int\varphi_{k}^{*}(x_{1})\;\varphi_{p}^{*}(x_{2})\;\rho_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;\varphi_{k}(x_{1}^{\prime})\;\varphi_{p}(x_{2}^{\prime})\;dx_{1}dx_{2}dx_{1}^{\prime}dx_{2}^{\prime}\;.$ (21)
The trace over space (18) is
${\rm Tr}\;\hat{\rho}_{2}\;=\;\sum_{kp}\langle\;kp\;|\;\hat{\rho}_{2}\;|\;kp\;\rangle\;=\;\int\rho_{2}(x_{1},x_{2},x_{1},x_{2})\;dx_{1}dx_{2}\;.$ (22)
Then the order index of the second-order density operator is defined as
$\omega(\hat{\rho}_{2})\;=\;\frac{\log\sup_{kp}N_{kp}}{\log{\rm Tr}\hat{\rho}_{2}}\;.$ (23)
In the same way, it is possible to define the $n$-th order reduced density operators
$\hat{\rho}_{n}\;=\;[\;\rho_{n}(x_{1},x_{2},\ldots,x_{n},x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime})\;]\;,$ (24)
with the matrix elements
$\rho_{n}(x_{1},x_{2},\ldots,x_{n},x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime})\;=$
$\langle\;\psi^{\dagger}(x_{n}^{\prime})\;\psi^{\dagger}(x_{n-1}^{\prime})\ldots\psi^{\dagger}(x_{1}^{\prime})\;\psi(x_{1})\;\psi(x_{2})\ldots\psi(x_{n})\;\rangle\;,$ (25)
acting on the Hilbert space
${\cal H}^{n}\;=\;{\cal H}_{1}\bigotimes{\cal H}_{1}\bigotimes{\cal H}_{1}\bigotimes\ldots\bigotimes{\cal H}_{1}\;.$ (26)
Then, it is straightforward to find the related order index $\omega(\hat{\rho}_{n})$. However, the reduced density matrices of order higher than two are rarely considered, since the main information is usually taken into account in the first-order and second-order reduced density matrices.
It is important to emphasize that the order indices for reduced density operators of any order are well defined for an arbitrary statistical system with a finite number of particles NN. This follows from the known properties of the reduced density matrices [22], for which the traces and norms of the related reduced density operators are finite for a finite number NN. This also directly follows from the fact that the density operators are semi-positive trace-class operators for which the inequality (6) is valid.
4 Correlation operators
Now we extend the notion of order indices to correlation operators that characterize a kind of order in a system and which can be treated as matrix elements of operators. For example, let us have a two-point correlation function
$C_{1}(x,x^{\prime})\;=\;\langle\;\hat{A}^{+}(x^{\prime})\;\hat{A}(x)\;\rangle$ (27)
composed of some local observables $\hat{A}(x)$. Then it can be considered as a matrix element of the first-order correlation operator
$\hat{C}_{1}\;=\;[\;C_{1}(x,x^{\prime})\;]$ (28)
acting on the Hilbert space $\mathcal{H}_{1}$ composed similarly to Ref. (9). The first-order index for the correlation operator (28) is given by
[…]
A second-order correlation operator
$\hat{C}_{2}\;=\;[\;C_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;]$ (30)
can be defined as being composed of the matrix elements
$C_{2}(x_{1},x_{2},x_{1}^{\prime},x_{2}^{\prime})\;=\;\langle\;\hat{A}^{+}(x_{2}^{\prime})\;\hat{A}(x_{1}^{\prime})\;\hat{A}(x_{1})\;\hat{A}(x_{2})\;\rangle\;.$ (31)
The second-order operator acts on the Hilbert space $\mathcal{H}^{2}$ similar to space (18). The order index for the correlation operator (30) is
[…]
The introduced order indices are the functions of the system parameters and size, so that varying these parameters and the number of particles it is possible to study the level of ordering in the system.
Here, for generality, we shall consider quantum systems. However, correlation functions can be defined for quantum as well as for classical systems. As soon as a correlation function $C_{n}(x_{1},x_{2},\ldots,x_{n},x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime})$ is given, independently of its origin, whether quantum or classical, the correlation operator $\hat{C}_{n}$ can be defined as above.
5 Bose-Einstein condensation
Let us illustrate the behavior of the order indices by concrete examples. We start with Bose-Einstein condensation that in the recent years has been widely studied both theoretically and experimentally, as can be inferred from review articles and books [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Here we study the dependence of order indices on the increasing system size.
The Hamiltonian of a system of spinless particles is
$\hat{H}\;=\;\int\psi^{\dagger}({\bf r})\;\left(-\;\frac{\nabla^{2}}{2m}\right)\;\psi({\bf r})\;d{\bf r}+\frac{1}{2}\int\psi^{\dagger}({\bf r})\;\psi^{\dagger}({\bf r}^{\prime})\;\Phi({\bf r}-{\bf r}^{\prime})\;\psi({\bf r}^{\prime})\;\psi({\bf r})\;d{\bf r}d{\bf r}^{\prime}\;,$ (33)
where $\Phi$ is the interaction potential. This potential, for a dilute system, can be taken in the local form
$\Phi({\bf r})\;=\;\Phi_{0}\;\delta({\bf r})\;,\qquad\Phi_{0}\;=\;4\pi\;\frac{a_{s}}{m}\;,$ (34)
in which $a_{s}$ is scattering length, assumed to be positive, and $m$, particle mass.
The possible appearance of Bose-Einstein condensate requires global gauge symmetry breaking, which is the necessary and sufficient condition for Bose condensation [30, 33, 37, 41], and which can be realized by the Bogolubov shift $\psi({\bf r})\;=\;\eta({\bf r})+\psi_{1}({\bf r})$, where
$\langle\;\psi_{1}({\bf r})\;\rangle\;=\;0\;,\qquad\int\eta^{*}({\bf r})\;\psi_{1}({\bf r})\;d{\bf r}\;=\;0\;.$ (35)
The number of condensed and non-condensed particles, respectively, are
$N_{0}\;=\;\int|\;\eta({\bf r})\;|^{2}d{\bf r}\;,\qquad N_{1}\;=\;\langle\;\hat{N}_{1}\;\rangle\;,$ (36)
with the operator of non-condensed particles
$\hat{N}_{1}\;=\;\int\psi_{1}^{\dagger}({\bf r})\;\psi_{1}({\bf r})\;d{\bf r}\;.$ (37)
The grand Hamiltonian, taking into account the normalizations (36), is
$H\;=\;\hat{H}-\mu_{0}N_{0}-\mu_{1}\hat{N}_{1}\;.$ (38)
Under the Bogolubov shift, there appear the expressions
$\rho_{1}({\bf r},{\bf r}^{\prime})\;\equiv\;\langle\;\psi_{1}^{\dagger}({\bf r}^{\prime})\;\psi_{1}({\bf r})\;\rangle\;,\qquad\sigma_{1}({\bf r},{\bf r}^{\prime})\;\equiv\;\langle\;\psi_{1}({\bf r}^{\prime})\;\psi_{1}({\bf r})\;\rangle\;,$ (39)
called the normal and anomalous averages, respectively. The first-order reduced density matrix becomes
$\rho({\bf r},{\bf r}^{\prime})\;=\;\eta^{*}({\bf r}^{\prime})\;\eta({\bf r})+\rho_{1}({\bf r},{\bf r}^{\prime})\;.$ (40)
For a uniform system, one has
$\eta\;=\;\sqrt{\rho_{0}}\;,\qquad\rho_{0}\;\equiv\;\frac{N_{0}}{V}\;.$ (41)
The operators of non-condensed particles can be expanded in plane waves,
$\psi_{1}({\bf r})\;=\;\sum_{k\neq 0}a_{k}\;\varphi_{k}({\bf r})\;,\qquad\varphi_{k}({\bf r})\;=\;\frac{1}{\sqrt{V}}\;e^{i{\bf k}\cdot{\bf r}}\;.$ (42)
The normal and anomalous averages take the form
$\rho_{1}({\bf r},{\bf r}^{\prime})\;=\;\frac{1}{V}\sum_{k\neq 0}n_{k}\;e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}\;,\qquad\sigma_{1}({\bf r},{\bf r}^{\prime})\;=\;\frac{1}{V}\sum_{k\neq 0}\sigma_{k}\;e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}\;,$ (43)
where
$n_{k}\;=\;\langle\;a_{k}^{+}\;a_{k}\;\rangle\;,\qquad\sigma_{k}\;=\;\langle\;a_{-k}\;a_{k}\;\rangle\;.$ (44)
We employ the self-consistent mean-field approach that is gapless and conserving [43, 44, 45, 46], and gives good quantitative agreement with Monte Carlo simulations for uniform Bose gas [47, 48, 49] and with Monte Carlo results [50, 51] for nonuniform trapped gas [52]. In this approximation, we have
$n_{k}\;=\;\frac{\omega_{k}}{2\varepsilon_{k}}\;\coth\left(\frac{\varepsilon_{k}}{2T}\right)-\;\frac{1}{2}\;,\qquad\sigma_{k}\;=\;-\;\frac{mc^{2}}{2\varepsilon_{k}}\;\coth\left(\frac{\varepsilon_{k}}{2T}\right)\;,$ (45)
where $T$ is temperature, the notation
$\omega_{k}\;=\;mc^{2}+\frac{k^{2}}{2m}$ (46)
is used, and the spectrum of collective excitations is
$\varepsilon_{k}\;=\;\sqrt{(ck)^{2}+\left(\frac{k^{2}}{2m}\right)^{2}}\;.$ (47)
The sound velocity $c$ is defined by the equation
$mc^{2}\;=\;\Phi_{0}(\rho_{0}+\sigma_{1})\;.$ (48)
The density of non-condensed particles and the summary anomalous average are
$\rho_{1}\;=\;\frac{1}{V}\sum_{k\neq 0}n_{k}\;,\qquad\sigma_{1}\;=\;\frac{1}{V}\sum_{k\neq 0}\sigma_{k}\;,$ (49)
the total particle density being $\rho=\rho_{0}+\rho_{1}$.
Let us consider the case of zero temperature, $T=0$. Then
$n_{k}\;=\;\frac{\omega_{k}-\varepsilon_{k}}{2\varepsilon_{k}}\;,\qquad\sigma_{k}\;=\;-\;\frac{mc^{2}}{2\varepsilon_{k}}\qquad(T=0)\;.$ (50)
The density of non-condensed particles is
$\rho_{1}\;=\;\frac{(mc)^{3}}{3\pi^{2}}\qquad(T=0)\;.$ (51)
The anomalous average takes the form
$\sigma_{1}\;=\;-mc^{2}\int\frac{1}{2\varepsilon_{k}}\;\frac{d{\bf k}}{(2\pi)^{3}}\qquad(T=0)\;.$ (52)
It is convenient to use the dimensionless quantities introducing the particle fractions
$n_{0}\;\equiv\;\frac{\rho_{0}}{\rho}\;,\qquad n_{1}\;\equiv\;\frac{\rho_{1}}{\rho}\;,\qquad\sigma\;\equiv\;\frac{\sigma_{1}}{\rho}\;,$ (53)
and the dimensionless sound velocities
$s\;\equiv\;\frac{mc}{\rho^{1/3}}\;,\qquad s_{B}\;\equiv\;\frac{mc_{B}}{\rho^{1/3}}\qquad\left(c_{B}\equiv\sqrt{\frac{\rho}{m}\;\Phi_{0}}\right)\;.$ (54)
The anomalous average (52), because of the use of the contact potential, diverges and requires regularization. We employ the dimensional regularization [39, 27] that is exact under asymptotically weak interactions, and can be analytically continued to finite interaction strength [46], which gives
$\sigma\;=\;\frac{s_{B}^{3}}{\pi^{2}}\left(n_{0}+\frac{s_{B}^{3}}{\pi^{2}}\;\sqrt{n_{0}}\right)^{1/2}\;.$ (55)
The interaction strength is characterized by the gas parameter
$\gamma\;\equiv\;\rho^{1/3}\;a_{s}\;,$ (56)
so that $s_{B}^{2}=4\pi\gamma$.
Finally, we obtain the closed set of equations for the fractions of condensed, $n_{0}$, and non-condensed, $n_{1}$, particles,
$n_{0}\;=\;1-\;\frac{s^{3}}{3\pi^{2}}\;,\qquad n_{1}\;=\;\frac{s^{3}}{3\pi^{2}}\;,$ (57)
the anomalous average
$\sigma\;=\;\frac{8}{\sqrt{\pi}}\;\gamma^{3/2}\left(n_{0}+\frac{8}{\sqrt{\pi}}\;\gamma^{3/2}\;\sqrt{n_{0}}\right)^{1/2}\;,$ (58)
and the sound velocity squared,
$s^{2}\;=\;4\pi\gamma(n_{0}+\sigma)\;.$ (59)
These equations allow us to define the condensate fraction, the fraction of uncondensed particles, the anomalous average, and the sound velocity as functions of the gas parameter. Then the order indices, being the functions of these variables, can also be defined as functions of $\gamma$. In order to understand the influence of interactions on the order in the system, we can find the explicit behavior of the solutions to the system of equations (57), (58), and (59) for small gas parameters:
$n_{0}\;\simeq\;1-\;\frac{8}{3\sqrt{\pi}}\;\gamma^{3/2}-\;\frac{64}{3\pi}\;\gamma^{3}-\;\frac{256}{9\pi^{3/2}}\;\gamma^{9/2}\;,$
$\sigma\;\simeq\;\frac{8}{\sqrt{\pi}}\;\gamma^{3/2}+\frac{64}{3\pi}\;\gamma^{3}-\;\frac{1408}{9\pi^{3/2}}\;\gamma^{9/2}\;,$
$s\;\simeq\;\sqrt{4\pi\gamma}+\frac{16}{3}\;\gamma^{2}-\;\frac{64}{9\sqrt{\pi}}\;\gamma^{7/2}\qquad(\gamma\rightarrow 0)\;.$ (60)
For arbitrary gas parameters, the system of equations can be solved numerically, which is shown in Fig. 1.
Figure 1: Condensate fraction $n_{0}$ (dashed-dotted line), fraction of uncondensed particles $n_{1}$ (dotted line), anomalous average $\sigma$ (dashed line), and dimensionless sound velocity $s$ (solid line) as functions of the gas parameter $\gamma$, at zero temperature.
The order indices of reduced density operators are calculated as is described above. For the first-order density operator, the occupation numbers are
$N_{k}\;=\;N_{0}\;\delta_{k0}+n_{k}\;.$ (61)
For $n_{k}$ and $\sigma_{k}$ in the case of a finite system, there exists the minimal wave vector
$k_{min}\;=\;\frac{2\pi}{L}\;=\;\frac{2\pi}{V^{1/3}}\;=\;2\pi\;\left(\frac{\rho}{N}\right)^{1/3}\;.$ (62)
The norm of the first-order density operator reads as
$\hat{\rho}_{1}\;||\;=\;\sup\left\{\;n_{0}N;\;\frac{s}{4\pi}\;N^{1/3}\;\right\}\;,$ (63)
which results in the order index
$\omega(\hat{\rho}_{1})\;=\;\frac{\ln\;\sup\left\{\;n_{0}N;\;\frac{s}{4\pi}\;N^{1/3}\;\right\}}{\ln N}\;.$ (64)
It is evident that for a macroscopic system in the thermodynamic limit, the whole system is ordered,
$\omega(\hat{\rho}_{1})\;\simeq\;1\qquad(N\rightarrow\infty)\;.$ (65)
The order does not appear suddenly, but it grows with increasing system size. The order index (64) is calculated numerically using the solutions to the system of equations (57), (58), and (59). The behavior of the order index for different gas parameters is presented in Fig. 2 as a function of $\ln N$. The increase of order depends on the value of the gas parameter. Thus for $\gamma=0.1$ there is no order for small NN, where the order index is negative, and the order quickly grows with NN. For $\gamma=0.5$, there is no order before $N=9$, and the order index reaches $1/2$ close to $N=77$.
The interactions suppress the level of order diminishing the order index, which is clearly seen for small $\gamma$, where
$\omega(\hat{\rho}_{1})\;\simeq\;1-\;\frac{8}{3\sqrt{\pi}\;\ln N}\;\gamma^{3/2}-\frac{224}{9\pi\;\ln N}\;\gamma^{3}\qquad(\gamma\rightarrow 0)\;.$ (66)
Figure 2: Appearance of order at Bose-Einstein condensation under increasing system size. First order index $\omega(\hat{\rho}_{1})$ as a function of $\ln N$ for different gas parameters $\gamma=0.1$ (solid line), $\gamma=0.5$ (dashed line), and $\gamma=1$ (dashed-dotted line).
The second-order reduced density matrix, taking account of the Bogolubov shift, and using the Hartree-Fock-Bogolubov (HFB) approximation, becomes
$\rho_{2}({\bf r}_{1},{\bf r}_{2},{\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;=\;\rho_{0}^{2}\;+$
$\rho_{0}\;[\;\rho_{1}({\bf r}_{1},{\bf r}_{1}^{\prime})+\rho_{1}({\bf r}_{1},{\bf r}_{2}^{\prime})+\rho_{1}({\bf r}_{2},{\bf r}_{1}^{\prime})+\rho_{1}({\bf r}_{2},{\bf r}_{2}^{\prime})+\sigma_{1}({\bf r}_{1},{\bf r}_{2})+\sigma_{1}^{*}({\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;]+$
$\rho_{1}({\bf r}_{1},{\bf r}_{1}^{\prime})\;\rho_{1}({\bf r}_{2},{\bf r}_{2}^{\prime})+\rho_{1}({\bf r}_{1},{\bf r}_{2}^{\prime})\;\rho_{1}({\bf r}_{2},{\bf r}_{1}^{\prime})+\sigma_{1}^{*}({\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;\sigma_{1}({\bf r}_{1},{\bf r}_{2})\;.$ (67)
The occupation numbers (21) read as
$N_{kp}\;=\;N_{0}^{2}\;\delta_{k0}\;\delta_{p0}+N_{0}\;(n_{k}\;\delta_{p0}+n_{p}\;\delta_{k0})+n_{k}\;n_{p}+\left(n_{k}^{2}+|\;\sigma_{k}\;|^{2}\right)\;\delta_{kp}\;.$ (68)
Using for the terms $n_{k}$ and $|\sigma_{k}|$ expressions (50), with the minimal wave vector (62), we obtain the norm
$\hat{\rho}_{2}\;||\;=\;\sup\left\{n_{0}^{2}N^{2};\;\frac{n_{0}s}{4\pi}\;N^{4/3};\;\frac{3s^{2}}{16\pi^{2}}\;N^{2/3}\right\}\;.$ (69)
The second-order index reads as
$\omega(\hat{\rho}_{2})\;=\;\frac{\ln\;\sup\left\{n_{0}^{2}N^{2};\;\frac{1}{4\pi}\;n_{0}sN^{4/3};\;\frac{3}{16\pi^{2}}\;s^{2}N^{2/3}\right\}}{2\ln N}\;.$ (70)
In the thermodynamic limit, we have
$\omega(\hat{\rho}_{2})\;\simeq\;1\qquad(N\rightarrow\infty)\;,$ (71)
while the behavior of the second-order index for finite values of NN is shown in Fig. 3. Again, we see that the growth of the order index depends on the interaction strength. Strong interactions make the development of order, under the increasing system size, more difficult.
Figure 3: Appearance of order at Bose-Einstein condensation under increasing system size. Second order index $\omega(\hat{\rho}_{2})$ as a function of $\ln N$ for different gas parameters $\gamma=0.1$ (solid line), $\gamma=0.5$ (dashed line), and $\gamma=1$ (dashed-dotted line).
6 Superconducting transition
Starting from the general Hamiltonian with attractive interactions,
$H\;=\;\sum_{s}\int\psi_{s}^{\dagger}({\bf r})\;\left(-\;\frac{\nabla^{2}}{2m}-\mu\right)\;\psi_{s}({\bf r})\;d{\bf r}\;+$
$\frac{1}{2}\sum_{ss^{\prime}}\int\psi_{s}^{\dagger}({\bf r})\;\psi_{s^{\prime}}^{\dagger}({\bf r}^{\prime})\;\Phi({\bf r}-{\bf r}^{\prime})\;\psi_{s^{\prime}}({\bf r}^{\prime})\;\psi_{s}({\bf r})\;d{\bf r}d{\bf r}^{\prime}\;,$ (72)
where the index $s$ labels spin variables, we pass to the momentum representation, assuming a uniform system, hence expanding the field operators over the plane waves,
$\psi_{s}({\bf r})\;=\;\frac{1}{\sqrt{V}}\sum_{k}c_{sk}\;e^{i{\bf k}\cdot{\bf r}}\;.$ (73)
Employing the Bardeen-Cooper-Schrieffer (BCS) restriction [53]
$c_{sk}\;c_{s^{\prime}p}\;=\;\delta_{-ss^{\prime}}\;\delta_{-kp}\;c_{sk}\;c_{-s,-k}\;,$ (74)
one comes to the BCS Hamiltonian
$H\;=\;\sum_{s}\;\sum_{k}\left(\frac{k^{2}}{2m}-\mu\right)\;c_{sk}^{\dagger}\;c_{sk}+\frac{1}{2V}\sum_{s}\;\sum_{kp}\Phi_{k-p}\;c_{sk}^{\dagger}\;c_{-s,-k}^{\dagger}\;c_{-s,-p}\;c_{sp}\;.$ (75)
For the interaction term, the Hartree-Fock-Bogolubov approximation yields
$c_{sk}^{\dagger}\;c_{-s,-k}^{\dagger}\;c_{-s,-p}\;c_{sp}\;=\;\delta_{kp}\left(n_{k}\;c_{sk}^{\dagger}\;c_{sk}+n_{k}\;c_{-s,-k}^{\dagger}\;c_{-s,-k}-n_{k}^{2}\right)\;+$
$\sigma_{p}\;c_{sk}^{\dagger}\;c_{-s,-k}^{\dagger}+\sigma_{k}\;c_{-s,-p}\;c_{sp}-\sigma_{k}^{*}\;\sigma_{p}\;,$ (76)
where the spin and momentum conservation in the binary statistical averages are taken into account,
$\langle\;c_{sk}^{\dagger}\;c_{s^{\prime}p}\;\rangle\;=\;\delta_{ss^{\prime}}\;\delta_{kp}\;n_{k}\;,\qquad n_{k}\;\equiv\;\langle\;c_{sk}^{\dagger}\;c_{sk}\;\rangle\;,$
$\langle\;c_{s^{\prime}p}\;c_{sk}\;\rangle\;=\;\delta_{-ss^{\prime}}\;\delta_{-kp}\;\sigma_{k}\;,\qquad\sigma_{k}\;\equiv\;\langle\;c_{-s,-k}\;c_{sk}\;\rangle\;.$ (77)
Thus we get the Hamiltonian
$H\;=\;\sum_{s}\;\sum_{k}\left(\frac{k^{2}}{2m}+\overline{\Phi}\;n_{k}-\mu\right)\;c_{sk}^{\dagger}\;c_{sk}+$
$\frac{1}{2V}\sum_{s}\;\sum_{kp}\Phi_{k-p}\left(\sigma_{p}\;c_{sk}^{\dagger}\;c_{-s,-k}+\sigma_{k}^{*}\;c_{-s,-p}\;c_{sp}\right)-$
$\frac{1}{2V}\sum_{s}\;\sum_{kp}\Phi_{k-p}\left(\sigma_{k}^{*}\;\sigma_{p}+\delta_{kp}\;n_{k}^{2}\right)\;,$ (78)
where
$\overline{\Phi}\;\equiv\;\frac{1}{V}\int\Phi({\bf r})\;d{\bf r}\;.$ (79)
To diagonalize the Hamiltonian, one resorts to the Bogolubov canonical transformation
$c_{sk}\;=\;u_{k}\;b_{sk}+v_{k}\;b_{-s,-k}^{\dagger}\;,$ (80)
in which
$|\;u_{k}\;|^{2}+|\;v_{k}\;|^{2}\;=\;1\;,\qquad u_{k}\;v_{-k}+u_{-k}\;v_{k}\;=\;0\;,$
$|\;u_{k}\;|^{2}\;=\;\frac{1}{2}\;\left(1+\frac{\omega_{k}}{\varepsilon_{k}}\right)\;,\qquad|\;v_{k}\;|^{2}\;=\;\frac{1}{2}\;\left(1-\;\frac{\omega_{k}}{\varepsilon_{k}}\right)\;.$
Here the notation
$\omega_{k}\;\equiv\;\frac{k^{2}}{2m}+\overline{\Phi}\;n_{k}-\mu$ (81)
is used. The spectrum of collective excitations has the form
$\varepsilon_{k}\;=\;\sqrt{\Delta_{k}^{2}+\omega_{k}^{2}}\;,$ (82)
where the gap is
$\Delta_{k}\;=\;-\;\frac{1}{2V}\sum_{p}\Phi_{k-p}\;\sigma_{p}\;.$ (83)
Then we get the diagonal Hamiltonian
$H\;=\;\sum_{s}\;\sum_{k}\varepsilon_{k}\;b_{k}^{\dagger}\;b_{k}+\sum_{k}(\omega_{k}-\varepsilon_{k}+\Delta_{k}\;\sigma_{k})\;,$ (84)
for which
$\langle\;b_{sk}^{\dagger}\;b_{sk}\;\rangle\;=\;\frac{1}{\exp(\varepsilon_{k}/T)-1}\;.$ (85)
The averages (77) become
$n_{k}\;=\;\frac{1}{2}-\;\frac{\omega_{k}}{2\varepsilon_{k}}\;\tanh\left(\frac{\varepsilon_{k}}{2T}\right)\;,\qquad\sigma_{k}\;=\;\frac{\Delta_{k}}{2\varepsilon_{k}}\;\tanh\left(\frac{\varepsilon_{k}}{2T}\right)\;.$ (86)
The chemical potential is defined by the equation
$\frac{2}{V}\sum_{k}n_{k}\;=\;2\int n_{k}\;\frac{d{\bf k}}{(2\pi)^{3}}\;=\;\rho\;.$ (87)
To find the first-order index, we need to consider the first-order reduced density operator
$\hat{\rho}_{1}\;=\;[\;\rho({\bf r},{\bf r}^{\prime})\;]\;,$ (88)
in which the density matrix is
$\rho({\bf r},{\bf r}^{\prime})\;=\;\langle\;\psi_{s}^{\dagger}({\bf r}^{\prime})\;\psi_{s}({\bf r})\;\rangle\;=\;\frac{1}{V}\sum_{k}n_{k}\;e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}\;.$ (89)
Then we have
$\hat{\rho}_{1}\;||\;=\;\sup_{k}n_{k}\;,\qquad{\rm Tr}\;\hat{\rho}_{1}\;=\;N\;.$ (90)
Hence the first-order index is
$\omega(\hat{\rho}_{1})\;=\;\frac{\ln\;\sup_{k}n_{k}}{\ln N}\;.$ (91)
Since $\sup n_{k}\leq 1/2$, we see that the first order index
$\omega(\hat{\rho}_{1})\;\simeq\;0\qquad(N\gg 1)$ (92)
does not show order.
The second-order reduced density operator
$\hat{\rho}_{2}\;=\;[\;\rho({\bf r}_{1},{\bf r}_{2},{\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;]$ (93)
is expressed through the matrix elements
$\rho({\bf r}_{1},{\bf r}_{2},{\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;=\;\langle\;\psi_{-s}^{\dagger}({\bf r}_{2}^{\prime})\;\psi_{s}^{\dagger}({\bf r}_{1}^{\prime})\;\psi_{s}({\bf r}_{1})\;\psi_{-s}({\bf r}_{2})\;\rangle\;,$ (94)
which in the HFB approximation read as
$\rho({\bf r}_{1},{\bf r}_{2},{\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;=\;\rho({\bf r}_{1},{\bf r}_{1}^{\prime})\;\rho({\bf r}_{2},{\bf r}_{2}^{\prime})+\sigma^{*}({\bf r}_{2}^{\prime},{\bf r}_{1}^{\prime})\;\sigma({\bf r}_{2},{\bf r}_{1})\;,$ (95)
where the anomalous average is
$\sigma({\bf r},{\bf r}^{\prime})\;=\;\langle\;\psi_{-s}({\bf r}^{\prime})\;\psi_{s}({\bf r})\;\rangle\;=\;\frac{1}{V}\sum_{k}\sigma_{k}\;e^{i{\bf k}\cdot({\bf r}-{\bf r}^{\prime})}\;.$ (96)
The anomalous average defines the number of pairs of correlated particles by the expression
$N_{\pi}\;=\;\int|\;\sigma({\bf r}_{2},{\bf r}_{1})\;|^{2}d{\bf r}_{1}d{\bf r}_{2}\;=\;\sum_{k}|\;\sigma_{k}\;|^{2}\;.$ (97)
Hence the fraction of pair-correlated particles is
$n_{cor}\;=\;\frac{2N_{\pi}}{N}\;.$ (98)
The trace of the density operator (93) takes the form
${\rm Tr}\;\hat{\rho}_{2}\;=\;\int\rho({\bf r}_{1},{\bf r}_{2},{\bf r}_{1},{\bf r}_{2})\;d{\bf r}_{1}d{\bf r}_{2}\;=\;N^{2}+\frac{1}{2}\;n_{cor}\;N\;.$ (99)
In the presence of the pair-correlated particles, the eigenvalue problem
$\int\rho({\bf r}_{1},{\bf r}_{2},{\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;\varphi({\bf r}_{1}^{\prime},{\bf r}_{2}^{\prime})\;d{\bf r}_{1}d{\bf r}_{2}\;=\;\lambda\;\varphi({\bf r}_{1},{\bf r}_{2})$ (100)
gives the normalized eigenfunction
$\varphi({\bf r}_{1},{\bf r}_{2})\;=\;\frac{\sigma({\bf r}_{2},{\bf r}_{1})}{\sqrt{N_{\pi}}}$ (101)
and the eigenvalue
$\lambda\;=\;\frac{1}{N_{\pi}}\sum_{k}n_{k}^{2}\;|\;\sigma_{k}\;|^{2}+N_{\pi}\;.$ (102)
As is seen from expressions (86), both $n_{k}$ and $|\sigma_{k}|$ quickly diminish with increasing $k$, so that
$\sum_{k}n_{k}^{2}\;|\;\sigma_{k}\;|^{2}\;\cong\;\sup_{k}n_{k}^{2}\sum_{k}|\;\sigma_{k}\;|^{2}\;.$
Then Eq. (102) can be written as
$\lambda\;=\;\sup_{k}n_{k}^{2}+N_{\pi}\;.$ (103)
Taking into account that $\sup_{k}n_{k}=1/2$, we have
$\hat{\rho}_{2}\;||\;=\;\frac{1}{4}+\frac{1}{2}\;n_{cor}\;N\;.$ (104)
Thus we come to the second order index
$\omega(\hat{\rho}_{2})\;=\;\frac{\ln\left(\frac{1}{4}+\frac{1}{2}\;n_{cor}N\right)}{\ln\left(N^{2}+\frac{1}{2}\;n_{cor}N\right)}\;.$ (105)
This shows that in thermodynamic limit
$\omega(\hat{\rho}_{2})\;\simeq\;\frac{1}{2}\qquad(N\rightarrow\infty)\;,$ (106)
provided $n_{cor}>0$. The behavior of the order index (105) as a function of arbitrary $\ln N$ is presented in Fig. 4. The order grows fast with the increase of the system size. Even for a small fraction of correlated particles $n_{cor}=0.1$, the order index becomes nonzero for $N>15$ and reaches $1/4$ for $N=404$.
Figure 4: Appearance of order at superconduction transition under increasing system size. Second order index $\omega(\hat{\rho}_{2})$ as a function of $\ln N$ for different fractions of pair-correlated particles $n_{cor}=0.1$ (solid line), $n_{cor}=0.5$ (dashed line), and $n_{cor}=1$ (dashed-dotted line).
7 Magnetic transition
To describe how magnetic order grows with the increasing system size, it is necessary to consider the correlation operators composed of spin operators. Let us consider spin operators $S_{i}^{z}$ corresponding to spin one-half ($S=1/2$). The matrix elements of the first-order correlation operator
$\hat{C}_{1}\;=\;[\;C_{ij}\;]$ (107)
are given by the correlation functions
$C_{ij}\;=\;\langle\;S_{i}^{z}\;S_{j}^{z}\;\rangle\;.$ (108)
The operator (107) acts on the Hilbert space
${\cal H}_{1}\;=\;\overline{{\cal L}}\left\{\varphi_{k}({\bf a}_{j})=\frac{1}{\sqrt{N}}\;e^{i{\bf k}\cdot{\bf a}_{j}}\right\}$ (109)
that is a closed linear envelope over a complete orthonormal basis, such that
$\sum_{i=1}^{N}\varphi_{k}^{*}({\bf a}_{i})\;\varphi_{p}({\bf a}_{i})\;=\;\delta_{kp}\;,\qquad\sum_{k}\varphi_{k}^{*}({\bf a}_{i})\;\varphi_{k}({\bf a}_{j})\;=\;\delta_{ij}\;.$
For concreteness, we consider spin $S=1/2$. Keeping in mind long-range interactions between the spins, we use the mean-field approximation for different lattice sites,
$\langle\;S_{i}^{z}\;S_{j}^{z}\;\rangle\;=\;\langle\;S_{i}^{z}\;\rangle\langle\;S_{j}^{z}\;\rangle\qquad(i\neq j)\;.$ (110)
In that way, the correlation function is
$\langle\;S_{i}^{z}\;S_{j}^{z}\;\rangle\;=\;S^{2}\;[\;\delta_{ij}+(1-\delta_{ij})\;M^{2}\;]\;,$ (111)
where the notation
$M\;\equiv\;\frac{1}{S}\;\langle\;S_{i}^{z}\;\rangle$ (112)
is used.
Employing the matrix element
$\langle\;k\;|\;\hat{C}_{1}\;|\;p\;\rangle\;=\;\sum_{ij}\varphi_{k}^{*}({\bf a}_{i})\;C_{ij}\;\varphi_{p}({\bf a}_{j})\;=\;S^{2}\;[\;1+\delta_{k0}\;\delta_{p0}\;M^{2}(N-1)\;]\;,$ (113)
we find the norm
$\hat{C}_{1}\;||\;=\;\sup_{k}\;\langle\;k\;|\;\hat{C}_{1}\;|\;k\;\rangle\;=\;S^{2}\;[\;1+M^{2}(N-1)\;]$ (114)
and the trace
${\rm Tr}\;\hat{C}_{1}\;=\;\sum_{i=1}^{N}C_{ii}\;=\;S^{2}N\;.$ (115)
Thus the first-order index is
$\omega(\hat{C}_{1})\;=\;\frac{\ln\{S^{2}[\;1+M^{2}(N-1)\;]\}}{\ln(S^{2}N)}\;.$ (116)
In the thermodynamic limit,
$\omega(\hat{C}_{1})\;\simeq\;1\qquad(N\rightarrow\infty)\;,$ (117)
as soon as $M>0$. The behavior of the order index (116) for arbitrary NN is shown in Fig. 5.
Figure 5: Appearance of order at magnetic transition under increasing system size. First order index $\omega(\hat{C}_{1})$ as a function of $\ln N$ for different magnetizations $M\equiv(1/S)\langle S_{i}^{z}\rangle$, with $M=0.01$ (solid line), $M=0.1$ (dashed line), and $M=0.5$ (dashed-dotted line).
The second-order correlation operator
$\hat{C}_{2}\;=\;[\;C_{ijmn}\;]$ (118)
is composed of the correlation functions
$C_{ijmn}\;\equiv\;\langle\;S_{i}^{z}\;S_{j}^{z}\;S_{m}^{z}\;S_{n}^{z}\;\rangle\;.$ (119)
Invoking the same procedure as above, with the use of the binary Ter Haar decoupling, we have
$C_{ijmn}\;=\;S^{4}\;\left\{\delta_{ij}\;\delta_{mn}+\delta_{ij}(1-\delta_{mn})\;M^{2}+(1-\delta_{ij})\;\delta_{mn}\;M^{2}+\right$.
$\left.+(1-\delta_{ij})\;(1-\delta_{mn})\;M^{4}\right\}\;.$ (120)
With the matrix element
$\langle\;kp\;|\;\hat{C}_{2}\;|\;kp\;\rangle\;=\;S^{4}\left\{\delta_{-kp}+2(\delta_{k0}\;\delta_{p0}\;N-\delta_{-kp})\;M^{2}\right$.+
$\left.(\delta_{k0}\;\delta_{p0}\;N^{2}-2\delta_{k0}\;\delta_{p0}\;N+\delta_{-kp})\;M^{4}\right\}\;,$ (121)
we obtain the norm
$\hat{C}_{2}\;||\;=\;\sup_{kp}\langle\;kp\;|\;\hat{C}_{2}\;|\;kp\;\rangle\;=\;S^{4}\;[\;1+(N-1)M^{2}\;]^{2}\;.$ (122)
The trace of the correlation operator (118) is
${\rm Tr}\;\hat{C}_{2}\;=\;\sum_{ij}C_{ijij}\;=\;S^{4}N^{2}\;.$ (123)
Therefore the second-order index coincides with the first,
$\omega(\hat{C}_{2})\;=\;\omega(\hat{C}_{1})\;.$ (124)
8 Crystal-liquid transition
The notion of order indices can be used for describing different types of arising order, whether off-diagonal or diagonal. As an illustration, let us show how the crystal order, arising under crystallization process can be treated. The starting point is the definition of appropriate correlation operators.
Let us consider the density operator
$\hat{\rho}({\bf r})\;\equiv\;\psi^{\dagger}({\bf r})\;\psi({\bf r})\;,$ (125)
whose statistical average gives the local particle density in the real space,
$\rho({\bf r})\;=\;\langle\;\hat{\rho}({\bf r})\;\rangle\;.$ (126)
The average density over the system is
$\rho\;=\;\frac{1}{V}\int\rho({\bf r})\;d{\bf r}\;=\;\frac{N}{V}\;.$ (127)
Let us label the locations of the maximal density by the vectors ${\bf a}_{j}$, with $j=1,2,\ldots,N$, so that
$\max_{{\bf r}}\rho({\bf r})\;=\;\rho({\bf a}_{j})\;.$ (128)
The fraction operator, associated with the points of the maxima, is given by the expression
$\hat{n}_{j}\;\equiv\;\frac{\hat{\rho}({\bf a}_{j})}{\rho}\;.$ (129)
And let us define the deviation operator
$\hat{D}_{j}\;\equiv\;\hat{n}_{j}-1$ (130)
characterizing the deviations of the local maximal density from the average density, so that
$D\;\equiv\;\langle\;\hat{D}_{j}\;\rangle\;=\;\langle\;\hat{n}_{j}\;\rangle-1\;=\;\langle\;\frac{\hat{\rho}({\bf a}_{j})}{\rho}\;\rangle-1\;.$ (131)
Assuming long-range interactions allowing for the mean-field approximation
$\langle\;\hat{n}_{i}\;\hat{n}_{j}\;\rangle\;=\;\langle\;\hat{n}_{i}\;\rangle\langle\;\hat{n}_{j}\;\rangle\qquad(i\neq j)\;,$ (132)
we have
$\langle\;\hat{D}_{i}\;\hat{D}_{j}\;\rangle\;=\;\langle\;\hat{D}_{i}\;\rangle\langle\;\hat{D}_{j}\;\rangle\;\equiv\;D^{2}\qquad(i\neq j)\;.$ (133)
For a liquid, with a uniform density, $D=0$, while for a crystal, whose density is nonuniform, $D>0$.
We define the correlation operator
$\hat{C}_{1}\;=\;[\;C_{ij}\;]\;,\qquad C_{ij}\;\equiv\;\langle\;\hat{D}_{i}\;\hat{D}_{j}\;\rangle\;,$ (134)
whose matrix elements are the correlation functions corresponding to the deviation operators. This operator acts on the Hilbert space (109).
With the matrix elements
$\langle\;k\;|\;\hat{C}_{1}\;|\;k\;\rangle\;=\;\langle\;\hat{D}_{j}^{2}\;\rangle+\delta_{k0}\;D^{2}(N-1)\;,$ (135)
we find the norm
$\hat{C}_{1}\;||\;=\;\langle\;\hat{D}_{j}^{2}\;\rangle+D^{2}(N-1)$ (136)
and the trace
${\rm Tr}\;\hat{C}_{1}\;=\;\langle\;\hat{D}_{j}^{2}\;\rangle\;N\;.$ (137)
Then we obtain the first-order index
$\omega(\hat{C}_{1})\;=\;\frac{\ln[\;\langle\;\hat{D}_{j}^{2}\;\rangle+D^{2}(N-1)\;]}{\ln(\langle\;\hat{D}_{j}^{2}\;\rangle N)}$ (138)
that is similar to the index (116) characterizing the arising magnetic order. In the thermodynamic limit,
$\omega(\hat{C}_{1})\;\simeq\;1\qquad(N\rightarrow\infty)\;,$ (139)
provided that $D>0$. The overall behavior of the order index (138) with respect to the varying particle number NN is analogous to that of the magnetic order index (116).
9 Conclusion
The process describing how the order grows with the increasing system size on the way to the thermodynamic limit is analyzed. For this purpose, a sequence of systems, with fixed parameters, but increasing system size, such that the number of particles NN and volume $V$ increase, while the density $\rho=N/V$ kept fixed, is considered. Each system in the sequence is assumed to be an equilibrium statistical system, which requires that the number of its constituents be sufficiently large, $N\gg 1$, and the system lifetime be sufficiently long, longer than the observation time. In this setup, no nonequilibrium effects need to be considered.
The arising and growing order, increasing together with the system size, can be quantified by the order indices that can be introduced for arbitrary trace-class operators, such as the reduced density operators, or other correlation operators. Of course, for quite small systems there could exist finite-size corrections. However, in small systems, where the order has not yet been developed, the order index is also quite small. Noticeable order emerges in sufficiently large systems with the number of particles $N\gg 1$, when boundary corrections do not play important role, and practically disappear for very large NN, completely vanishing in the thermodynamic limit when $N\rightarrow\infty$.
With the increasing system size, hence the increasing number of particles NN, the order index grows from zero or even negative, showing the absence of order, to finite positive values, reaching, in the thermodynamic limit, $N\rightarrow\infty$ the maximal possible value. Thus, for the cases where the order is due to some arrangement of individual variables, the order index reaches the value one, while in the case, where the order is caused by pairing correlations, the single-particle order index shows no order and the second-order index can reach the value $1/2$. The order indices, increasing with the growing system size, are considered for the processes of Bose-Einstein condensation, superconductivity, magnetization, and crystallization.
In conclusion, it is useful to emphasize once more why the introduced notion of order indices can describe the emergence and growth of order under the increasing system size and to remind the qualitative physical picture behind the formulas.
In a microscopic system of just a few particles there can be no order. When the system increases reaching mesoscopic size, where the number of particles NN is sufficiently large, there can appear a metastable ordered phase having a finite lifetime. More detailed discussion of characteristic times is postponed to the Appendix. Here we concentrate on spatial correlations that are in the origin of arising order.
Correlation length $l_{cor}$ is connected with the reduced density matrix $\rho({\bf r},{\bf r}^{\prime})$ and can be defined by the relation
$\rho l_{cor}^{3}\;=\;\sup_{k}\int\varphi_{k}^{*}({\bf r})\;\rho({\bf r},{\bf r}^{\prime})\;\varphi_{k}({\bf r}^{\prime})\;d{\bf r}d{\bf r}^{\prime}\;,$ (140)
thus specifying the volume where the density matrix is essentially nonzero. By using the notation for the average distance between particles, $a$, given by the identity
$\rho a^{3}\;=\;1\;,$ (141)
and noticing that the right-hand side of (140) is the norm of $\hat{\rho}_{1}$, we have
[…]
In order to show that these equations really define correlation length, let us consider the limiting cases. When the density matrix describes the ultimate short-range correlations, so that $\rho({\bf r},{\bf r}^{\prime})\sim\delta({\bf r}-{\bf r}^{\prime})$, then $l_{cor}\sim a$. In the opposite case, when there exist long-range correlations, such that $\rho({\bf r},{\bf r}^{\prime})\sim\rho$ throughout the whole system of length $L\sim V^{1/3}$, then (140) gives $l_{cor}\sim L$. This behavior perfectly corresponds to physical intuition that the system is completely ordered when the correlation length is comparable to the system length. In this way, the increase of the density operator norm leads to the increase of the correlation length and to the rise of the order index.
The correlation length $l_{cor}$ characterizes correlations between single particles. It may happen that single particles do not exhibit long-range correlations, but pairs of particles do experience long-range correlations, as it happens for superconductors. This implies that there exist, actually, different correlation lengths, with respect to single particles and pairs of particles. The correlation length for pairs of particles is defined by the equation
[…]
This correlation length can vary between the limiting cases of the average particle distance $a$ and the system length $L$.
In the same way, it is straightforward to introduce the correlation length $l_{cor}^{(n)}$ characterizing correlations between the groups of $n$ particles,
[…]
Thence, the growth of a correlation length implies the increase of the operator norm and of the corresponding order index. In a similar way, correlation lengths can be defined for correlation operators of different nature, e.g., based on correlation functions for spin operators,
[…]
As has been noticed, correlations and, respectively, the ordering can be absent for single particles but present for particle pairs, as in superconductors. General relations between the existence of ordering in correlation operators depend on particular systems and concrete types of studied ordering. Below, we summarize our knowledge on the existence of ordering and the behavior of order indices for correlation operators. For concreteness, we can keep in mind the reduced density operators $\hat{\rho}_{n}$ that are a particular kind of correlation operators.
Conjecture 1. If some kind of ordering happens in a system, there should exist a related correlation operator and a corresponding order index depending on the system size.
Conjecture 2. If there is a long-range order in $\hat{\rho}_{1}$, it also exists in all higher orders of $\hat{\rho}_{n}$, with admissible $n=2,3,\ldots$.
Conjecture 3. If there is no order in $\hat{\rho}_{1}$, this does not mean its absence in some higher orders of $\hat{\rho}_{n}$.
Conjecture 4. If there is long-range order in $\hat{\rho}_{2}$, it exists in all even-orders of $\hat{\rho}_{2n}$, with admissible $n=2,3,\ldots$.
Effects of asymptotic symmetry breaking can be observed in different finite systems, such as trapped atoms, quantum dots, metallic grains, and atomic nuclei [11]. The notion of order indices can be useful for the description of and manipulation with these finite quantum systems.
Appendix. Characteristic times
To better understand the physics of processes occurring in finite statistical systems, and the meaning of equilibrium in such systems, it is worth recalling several important facts.
The notion of thermodynamic equilibrium presupposes that the system can be observed for a sufficiently long time, longer than the time of local fluctuations, so that what is observed and measured in an equilibrium system are its average properties [1, 2, 3, 4, 5]. In an infinite system, an equilibrium thermodynamics phase, under appropriate conditions, can live infinite time. In a finite system, there are restrictions on the lifetime of a metastable state.
First of all, there is a restriction from below, requiring that the lifetime be much longer than the local equilibration time [4, 54, 55], otherwise no ordered phase can arise. This time can be estimated as $t_{loc}\sim\lambda/v$, where $\lambda\sim 1/(\rho r_{int}^{2})\sim a^{3}/r_{int}$ is the mean free path, $v\sim\hbar/(mr_{int})$, characteristic particle velocity, $a$, mean interparticle distance, $r_{int}$, interaction radius, and $m$, particle mass. Thus the local equilibration time is $t_{loc}\sim ma^{3}/{\hbar r}_{int}$. For the parameters typical of condensed matter, the local equilibration time is of the order $10^{-13}$ s.
In some cases, there can appear mesoscopic fluctuations involving several particles and living longer than the local equilibration time. The size of such fluctuations is much larger than the mean interparticle distance $a$ that in condensed matter is usually close to the interaction radius. The mesoscopic fluctuations representing one phase inside another are called heterophase fluctuations [54, 56, 57]. They can occur when their existence diminishes the system free energy. The typical lifetime of such fluctuations in condensed matter is $t_{mes}\sim 10^{-12}$ s. In the interval between $t_{loc}$ and $t_{mes}$, the system has yet to be treated as nonequilibrium, becoming equilibrium for the times much longer than $t_{mes}$.
Thus, in the situation, where there are no strongly nonequilibrium initial conditions, the system can be considered as equilibrium when the observation time is much longer then the local equilibration time $t_{loc}$ and the mesoscopic fluctuation time $t_{mes}$, while from above, the observation time is limited by the lifetime $t_{met}$ of the considered metastable state. In this way, a system can be handled as equilibrium if the observation time $t_{obs}$ is in the interval of times
$t_{loc}\ll t_{mes}\ll t_{obs}\ll t_{met}\;.$
The lifetime of a metastable state can be estimated in the following way. Let the space of microscopic states of the considered system be a Hilbert space $\mathcal{H}$. Different phases are characterized by different subspaces of the total system space. Say the subspace of typical disordered states is $\mathcal{H}_{0}$ and the subspace of typical ordered states is $\mathcal{H}_{1}$. The separation of the ordered and disordered states can be done by means of the method of restricted trace [54, 55, 58, 59, 60]. The subspaces $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ can be represented as weighted Hilbert spaces [61, 62], with the weights selecting the required typical states [54, 55].
Suppose the space $\mathcal{H}_{0}$ possesses a vacuum state $|0\rangle$, while the space $\mathcal{H}_{1}$ possesses a vacuum state $|0\rangle_{1}$. All other states with the corresponding typical properties of the studied phases are formed by generating them from the corresponding vacuum states. The lifetime of a metastable phase can be estimated as the escape time required for the transition from one phase to another [63]. The escape rate (or transition rate) reads as $\gamma_{esc}=p_{esc}/t_{0}$, where $p_{esc}$ is the escape probability $p_{esc}=|\langle 0|0\rangle_{1}|^{2}$ and $t_{0}$ is a characteristic time of the order of the local equilibration time. The rates are inversely proportional to the related times, so that $\gamma_{esc}\equiv 1/t_{esc}$. Thus the phase lifetime is given by the expression
$t_{esc}\;=\;\frac{t_{0}}{|\;\langle\;0\;|\;0\;\rangle_{1}\;|^{2}}\;.$ (146)
The transition between different phases, characterized by different Hilbert spaces, is somewhat similar, although mathematics is of course rather different, to the transition between different parts of the same Hilbert space defining the localization time [64]. The other analogy is the transition (escape) from one potential well into another well in real space [63].
To be more precise and to better understand how the expression for the escape time can be derived, let us illustrate this by a concrete example of a transition from the normal, uncondensed, phase to the Bose-condensed phase. Mathematical details describing systems with Bose-Einstein condensate can be found in Refs. [34, 65].
Let us consider a system of Bose particles characterized by the field operators $\psi({\bf r})$ satisfying Bose commutation relations. Strictly speaking, the field operator depends as well on time which we do not write explicitly for the sake of simplicity of notation. For the same reason, for a while we do not mention internal degrees of freedom, if any, which can be included by treating the field operators as columns with respect to these variables. We do not want to complicate the consideration, when it is not important, as far as our principal aim is to study the growth of order under the increase of the system size in real space. When necessary, we can include internal degrees of freedom, e.g. spin.
Defining the vacuum state, such that
$\psi({\bf r})\;|\;0\;\rangle\;=\;0\;,$ (147)
it is possible to generate the states
$|\;\varphi\;\rangle\;=\;\sum_{n=0}^{\infty}\frac{1}{\sqrt{n!}}\int\varphi({\bf r}_{1},{\bf r}_{2},\ldots,{\bf r}_{n})\;\prod_{i=1}^{n}\psi^{\dagger}({\bf r}_{i})\;d{\bf r}_{i}\;|\;0\;\rangle\;,$ (148)
where $\varphi({\bf r}_{1},{\bf r}_{2},\ldots,{\bf r}_{n})$ is a symmetric function of its arguments. The closed linear envelope of all states $|\varphi\rangle$ generated from the vacuum state $|0\rangle$ by the field operators $\psi$, forms the Fock space
${\cal F}(\psi)\;=\;\overline{{\cal L}}\{\;|\;\varphi\;\rangle\;\}\;.$ (149)
The standard system Hamiltonian, with pair interactions, is invariant with respect to the global gauge symmetry, when the operators $\psi({\bf r})$ are replaced by $\psi({\bf r})e^{i\alpha}$, with a real phase $\alpha$.
If in the system there starts appearing Bose-Einstein condensate, its correct description requires the occurrence of gauge symmetry breaking, since this breaking is the necessary and sufficient condition for Bose-Einstein condensation [30, 33, 37, 38, 39]. The most convenient way of gauge symmetry breaking is the use of the Bogolubov shift [1, 2] that is an exact canonical transformation
$\psi({\bf r})\;=\;\eta({\bf r})+\psi_{1}({\bf r})\;.$ (150)
Here the field operator $\psi_{1}$, satisfying the Bose commutation relations, is the field operator of uncondensed particles and $\eta$ is the condensate wave function normalized to the number of condensed particles
$N_{0}\;=\;\int|\;\eta({\bf r})\;|^{2}d{\bf r}\;.$ (151)
As is easy to notice, the state $|0\rangle$, which is a vacuum with respect to $\psi$, is not a vacuum with respect to $\psi_{1}$, since
$\psi_{1}({\bf r})\;|\;0\;\rangle\;=\;-\eta({\bf r})\;|\;0\;\rangle$ (152)
is not zero. Hence, with respect to $\psi_{1}$, there exists another vacuum defined by the condition
$\psi_{1}({\bf r})\;|\;0\;\rangle_{1}\;=\;0\;.$ (153)
From this vacuum, it is straightforward to generate the functions
$|\;\varphi_{1}\;\rangle\;=\;\sum_{n=0}^{\infty}\frac{1}{\sqrt{n!}}\int\varphi_{1}({\bf r}_{1},{\bf r}_{2},\ldots,{\bf r}_{n})\;\prod_{n=1}^{n}\psi_{1}^{\dagger}({\bf r})\;d{\bf r}_{i}\;|\;0\;\rangle_{1}$ (154)
forming the Fock space as their closed linear envelope
${\cal F}(\psi_{1})\;=\;\overline{{\cal L}}\{\;|\;\varphi_{1}\;\rangle\;\}\;.$ (155)
Canonical transformations can be realized by transformation operators [66]. For example, the Bogolubov shift (150) can be represented as due to the action of the Bogolubov operator
[…]
in which
$\hat{D}\;=\;\int[\;\eta^{*}({\bf r})\;\psi({\bf r})-\eta({\bf r})\;\psi^{\dagger}({\bf r})\;]\;d{\bf r}\;.$ (157)
The Bogolubov shift (150) is equivalent to the transformation
$\psi({\bf r})\;=\;\hat{B}\;\psi_{1}({\bf r})\;\hat{B}^{-1}\;,\qquad\psi_{1}({\bf r})\;=\;\hat{B}^{-1}\;\psi({\bf r})\;\hat{B}\;$. (158)
Respectively, the vacuum states are related by the transformations
$|\;0\;\rangle\;=\;\hat{B}\;|\;0\;\rangle_{1}\;,\qquad|\;0\;\rangle_{1}\;=\;\hat{B}^{-1}|\;0\;\rangle\;.$ (159)
For two operators $\hat{A}_{1}$ and $\hat{A}_{2}$, whose commutator $\hat{A}_{1},\hat{A}_{2}]$ commutes with both $\hat{A}_{1}$ and $\hat{A}_{2}$, the Baker-Hausdorff formula is valid,
$\exp(\hat{A}_{1}+\hat{A}_{2})\;=\;e^{\hat{A}_{1}}\;e^{\hat{A}_{2}}\;\exp\left(-\;\frac{1}{2}\;[\;\hat{A}_{1},\;\hat{A}_{2}\;]\right)\;.$
Using this, we have
$\hat{B}^{-1}\;=\;\exp\left\{\int\eta({\bf r})\;\psi^{\dagger}({\bf r})\;d{\bf r}\right\}\;\exp\left\{-\int\eta^{*}({\bf r})\;\psi({\bf r})\;d{\bf r}\right\}\;\exp\left\{-\;\frac{1}{2}\int|\;\eta({\bf r})\;|^{2}\;d{\bf r}\right\}\;.$ (160)
Acting by the operator (160) on the vacuum state $|0\rangle$, we find
$\hat{B}^{-1}|\;0\;\rangle\;=\;\exp\left\{-\;\frac{1}{2}\int|\;\eta({\bf r})\;|^{2}\;d{\bf r}\right\}\;\exp\left\{\int\eta({\bf r})\;\psi^{\dagger}({\bf r})\;d{\bf r}\right\}\;|\;0\;\rangle\;.$ (161)
Here it is useful to remember that a coherent state that is the solution of the equation
$\psi({\bf r})\;|\;\eta\;\rangle\;=\;\eta({\bf r})\;|\;\eta\;\rangle\;,$ (162)
where $\eta({\bf r})$ is a coherent field [38], has the form
$|\;\eta\;\rangle\;=\;\exp\left\{-\;\frac{1}{2}\int|\;\eta({\bf r})\;|^{2}\;d{\bf r}\right\}\;\exp\left\{\int\eta({\bf r})\;\psi^{\dagger}({\bf r})\;d{\bf r}\right\}\;|\;0\;\rangle\;$ (163)
coinciding with the right-hand side of Eq. (161). Hence the vacuum $|0\rangle_{1}$ is a normalized coherent state,
$|\;0\;\rangle_{1}\;=\;\hat{B}^{-1}\;|\;0\;\rangle\;=\;|\;\eta\;\rangle\;.$ (164)
This is in agreement with the equation
$\psi({\bf r})\;|\;0\;\rangle_{1}\;=\;\eta({\bf r})\;|\;0\;\rangle_{1}$ (165)
following from the form (150) of the Bogolubov shift.
Taking into account normalization (151), we obtain
 $\langle\;0\;|\;0\;\rangle_{1}\;=\;\exp\left(-\frac{1}{2}\;N_{0}\right)\;.$ (166)
Then the transition probability is
$|\;\langle\;0\;|\;0\;\rangle_{1}\;|^{2}\;=\;\exp(-N_{0})\;.$ (167)
Therefore the escape time (146) reads as
$t_{esc}\;=\;t_{0}\;e^{N_{0}}\;.$ (168)
Formula (168) shows that if the system is small and the ordered phase is also small, such that $N_{0}$ is not much larger than one, it is not stable and, occasionally appearing, disappears very quickly. But in a larger system that can be sufficiently large, it becomes more stable and can live longer. Thus, with the increasing system size, the possibility of a larger ordered phase increases, which implies the growing order. In the thermodynamic limit, when $N_{0}\propto N\rightarrow\infty$, an absolutely equilibrium ordered phase can become absolutely stable.
It is useful to mention that, in general, the phase factor of a coherent state is not defined, so that there can appear the condensate with the coherent field $\eta e^{i\alpha}$ that is normalized to the same number $N_{0}$ as in the integral (151), and with an arbitrary phase $\alpha$. The phase appears randomly. But after appearing, it is preserved as far as the transition to another condensate with a different phase $\alpha\neq 2\pi n$, with $n=0,1,2,\ldots$, is suppressed, since the transition rate between different condensates
$|\;\langle\;\eta\;|\;\eta\;e^{i\alpha}\;\rangle\;|^{2}\;=\;\left|\;\exp\left\{-N_{0}\left(1-e^{i\alpha}\right)\right\}\;\right|^{2}\;=\;\exp\left\{-4N_{0}\;\sin^{2}\left(\frac{\alpha}{2}\right)\right\}$
is close to zero for large $N_{0}$ and $\alpha\neq 2\pi n$. Therefore the first randomly appearing condensate phase prevails and, without the loss of generality, it can be set to zero.
The picture, delineated above for the appearing Bose condensate under the increase of the system size, remains similar for the arising and growing order of any kind. In a microscopic system, there is no yet order. In a mesoscopic system there arises the embryo of the ordered phase, which lives finite time $t_{esc}$. Then, under the appropriate conditions, with the growth of the system, the size of the ordered phase grows and is getting more and more stable. The escape time increases by the law
$t_{esc}\;=\;t_{0}\;e^{cN_{0}}\;,$ (169)
where $t_{0}$ is of the order of the local equilibration time, which is the minimal time for establishing local equilibrium required for the ordered phase appearance, $c$ is a constant of order one, and $N_{0}$ is the number of particles in the ordered phase. Expression (169), to some extent, reminds us the form of the Arrhenius law [67], however the meaning is rather different. The Arrhenius law defines the typical time (or rate) for the transition between locations in the real space, while here $t_{esc}$ defines the transition between Fock spaces.
The typical lifetime $t_{esc}$, is estimated by the formula (169), where $t_{0}$ is of the order of local equilibration time of $10^{-13}$ s and $c$ is a constant of order one, characterizing the system. As is seen, for the number of particles in the ordered phase $N_{0}$ of order one, the lifetime does not differ much from the local equilibration time, while for $N_{0}$ of order $10$, the lifetime can be of order $10^{-9}$ s. For the number of particles in the system of order $50$, the lifetime of the ordered phase reaches the order of $16$ years. With increasing NN, the lifetime quickly grows to huge values. This explains why phase transitions and ordered phases that, strictly speaking, in finite systems are metastable, do exist and can be observed without any problem.
Declarations
Author contributions: All authors contributed to the study conception and design. Material preparation, data collection and analysis were performed by V.I. Yukalov and E.P. Yukalova. The first draft of the manuscript was written by V.I. Yukalov and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Funding: No funding was received to assist with the preparation of this manuscript.
Financial interests: The authors have no relevant financial or non-financial interests to disclose.
Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.